OpenStudy (okdutchman7):

Please help. WILL MEDAL! @surjithayer @mathmale @ganeshie8 @Hero @TheSmartOne @sleepyjess @Preetha @SolomonZelman @zepdrix @rebeccaxhawaii @Vocaloid @dtan5457 @MrNood

OpenStudy (okdutchman7):

\[-2 + \sqrt{3} \sec \theta = 0\]

OpenStudy (mathmale):

We communicated about this problem earlier. What happened? Why repost it? I've suggested that you replace that "sec theta" temporarily with the variable z. How would you then solve for z? Please show your work.

OpenStudy (okdutchman7):

I'm clueless on how to do this type of problem.

OpenStudy (okdutchman7):

Can someone teach me how to do this?

OpenStudy (okdutchman7):

@mathmale You didn't actually help me solve it at all.

OpenStudy (fortytherapper):

As mathmale said \[-2 + \sqrt{3}*z = 0\] You wanted to find what sec(theta) is, so instead find what z is first by using algebra

OpenStudy (mathmale):

Please solve this equation for z.

OpenStudy (okdutchman7):

I don't understand.

OpenStudy (mathmale):

All right. Please add 2 to both sides of this equation. Your result?

OpenStudy (fortytherapper):

It's the same as finding z for something like: \[1+z=2\]

OpenStudy (okdutchman7):

\[\sqrt{3} \times z = 2\]

OpenStudy (mathmale):

Good. Now, divide both sides by sqrt(3) and you will then have solved for z.

OpenStudy (okdutchman7):

\[z = \frac{ 2 }{ sqrt{3} } \]

OpenStudy (mathmale):

Very good. Now, replace that 'z' with sec theta.

OpenStudy (okdutchman7):

z \[z = \frac{ 2\sqrt{3} }{ 3 }\]

OpenStudy (okdutchman7):

\[\sec \theta = \frac{ 2\sqrt{3} }{ 3 }\]

OpenStudy (mathmale):

Yes. Since you are probably more comfortable with the cosine function than you are with the secant function, try inverting both sides of the equation. sec theta becomes 1 / sec theta = cos theta. Can you finish this work?

OpenStudy (mathmale):

Left side: cos theta. Right side = ?

OpenStudy (okdutchman7):

The instructions tell me to solve the following equations given that \[0degrees \le \theta < 360 degrees\]

OpenStudy (mathmale):

Yes. We're getting there. what's the right side going to look like?

OpenStudy (fortytherapper):

\[\frac{ 2\sqrt{3} }{ 3 }\] Why this? I think I missed something

OpenStudy (okdutchman7):

I'm confused again.

OpenStudy (okdutchman7):

I got lost here: secθ=23√3

OpenStudy (mathmale):

Let me quickly summarize what we've done so far: \[-2 + \sqrt{3} \sec \theta = 0\]

OpenStudy (mathmale):

can be solved for sec theta: sec theta = 2/sqrt(3).

OpenStudy (mathmale):

\[\sec \theta = \frac{ 2 }{ \sqrt{3} }\]

OpenStudy (mathmale):

Agree or not?

OpenStudy (okdutchman7):

I understand

OpenStudy (mathmale):

Good. Now, invert both sides. Recall that sec theta = 1 / cos theta.

OpenStudy (okdutchman7):

\[\sec \theta = \frac{ 2\sqrt{3} }{ 3 }\]

OpenStudy (okdutchman7):

This is where I'm stuck.

OpenStudy (okdutchman7):

I don't know what to do next.

OpenStudy (mathmale):

"Invert" means "to turn something upside down." Inverting "sec theta" produces 1 over sec theta = 1 / sec theta and inverting 2 / sqrt(3) produces sqrt(3) / 2. Clear on this?

OpenStudy (okdutchman7):

I understand that cosine is the inverse function of secant

OpenStudy (mathmale):

Sorry to say "no" on that, but that's not the case; you meant " cosine is the reciprocal of the secant."

OpenStudy (okdutchman7):

oh yes good catch

OpenStudy (okdutchman7):

:3

OpenStudy (okdutchman7):

You understand what I mean.

OpenStudy (mathmale):

Inverting "sec theta" produces 1 over sec theta = 1 / sec theta and inverting 2 / sqrt(3) produces sqrt(3) / 2. Clear on this? then we g et 1 / sec theta = cos theta = sqrt(3) / 2:

OpenStudy (mathmale):

\[\cos \theta = \frac{ \sqrt{3} }{ 2 }\]

OpenStudy (okdutchman7):

So, \[ \frac{ 1 }{ \sec \theta } = \frac{ \sqrt{3} }{ 2 }\]

OpenStudy (mathmale):

When I see this latest equation, I think immediately: since cos theta = adjacent side / hyp, cos theta in this problem = sqrt(3) / 2, and adj side =sqrt(3), and hyp = 2. Clear, or need more discussion?

OpenStudy (okdutchman7):

Okay I understand.

OpenStudy (okdutchman7):

clear

OpenStudy (mathmale):

\[\frac{ 1 }{ \sec \theta } = \frac{ \sqrt{3} }{ 2 }=\cos \theta\]

OpenStudy (okdutchman7):

gottit

OpenStudy (mathmale):

Yes. You merely replace 1/sec theta with cos theta.

OpenStudy (okdutchman7):

understood

OpenStudy (mathmale):

so now we have:\[\cos \theta = \frac{ \sqrt(3) }{ 2 }\]

OpenStudy (okdutchman7):

Okay, I'm still with you.

OpenStudy (mathmale):

we can interpret this to mean that cos theta = adj. side / hypotenuse = sqrt(3) / 2. Thus, I claim that the hyp has a length of 3 and the adj side has a length of sqrt(3). OK?

OpenStudy (mathmale):

Sorry, hyp has length 2.

OpenStudy (okdutchman7):

Understood.

OpenStudy (mathmale):

OK. While not strictly necessary, sketching this angle might help you figure out what the angle theta is.

OpenStudy (okdutchman7):

So it is a 30-60-90 triangle.

OpenStudy (mathmale):



OpenStudy (mathmale):

Yes, that's correct. adj side is sqrt(3); hyp is 2. Which angle have we here (theta)? 30? 60? 90?

OpenStudy (okdutchman7):

So now I have to find all of the values between 0 and 360 degrees

OpenStudy (okdutchman7):

30

OpenStudy (mathmale):

You're saying that the cosine of 30 degrees is sqrt(3) / 2. True. If I weren't here to say, "True," how would you check that cos 30 actually equals sqart(3) / 2?

OpenStudy (okdutchman7):

How can I tell if there are any other values?

OpenStudy (okdutchman7):

I'm not sure how to check.

OpenStudy (okdutchman7):

Put 30 into the cosine function?

OpenStudy (mathmale):

Experimentation. 30 degrees , as measured from the positive x-axis, has an adj side of sqart(3) and a hyp of 2. What would happen if you were to graph -30 degrees? what would the adj side be in that case? what would the length of the hyp be in that case?

OpenStudy (okdutchman7):

The sides would be the same length at -30 degrees.

OpenStudy (mathmale):

If you were to put 30 deg into the cosine function, you'd just get sqrt(3)/2, or approx. 0.866.